# All-in-One Slider for Attribute Manipulation in Diffusion Models

### v1

@@ Abstract
Text-to-image (T2I) diffusion models have made significant strides in generating high-quality images. However, progressively manipulating certain attributes of generated images to meet the desired user expectations remains challenging, particularly for content with rich details, such as human faces. Some studies have attempted to address this by training slider modules. However, they follow a One-for-One manner, where an independent slider is trained for each attribute, requiring additional training whenever a new attribute is introduced. This not only results in parameter redundancy accumulated by sliders but also restricts the flexibility of practical applications and the scalability of attribute manipulation. To address this issue, we introduce the All-in-One Slider, a lightweight module that decomposes the text embedding space into sparse, semantically meaningful attribute directions. Once trained, it functions as a general-purpose slider, enabling interpretable and fine-grained continuous control over various attributes. Moreover, by recombining the learned directions, the All-in-One Slider supports zero-shot manipulation of unseen attributes (e.g., races and celebrities) and the composition of multiple attributes. Extensive experiments demonstrate that our method enables accurate and scalable attribute manipulation, achieving notable improvements compared to previous methods. Furthermore, our method can be extended to integrate with the inversion framework to perform attribute manipulation on real images, broadening its applicability to various real-world scenarios. The code and trained model will be released at: https://github.com/ywxsuperstar/KSAE-FaceSteer.
Introduction
@@
With the widespread emergence of diffusion models (Wu et al. 2023; Podell et al. 2023), text-to-image (T2I) generation has advanced significantly in recent years. Despite their impressive performance, a major challenge remains: providing users with continuous and fine-grained control functions over specific image attributes, which is essential for users to adjust generated images to obtain the desired results. As illustrated by the example in Figure 1(1), traditional prompt engineering approaches, such as simply appending phrases like “with a big smile” to the base prompt, “A photo of a gentle Thai man with layered hair, in a warm-toned studio”, often result in coarse-grained and rigid manipulations. Such textual variations not only prevent users from specifying the intensity of the smile but also inadvertently affect many unrelated attributes, such as the subject’s hairstyle, pose, and identity, leading to suboptimal outcomes.
@@
Figure 1: Our All-in-One Slider shows advantages in: (1) Fine-grained and continuous control over desired attribute, without affecting other attributes (e.g., subjects’ identity and appearance). (2) Combination of multiple facial attributes (e.g., smile and age) for consistent and conflict-free transformations. (3) Zero-shot generalization to unseen attributes, without multiple and cumbersome training processes.
@@
To enable attribute manipulation, early studies focused on exploring the representation properties of generative models and identifying the semantic directions corresponding to attributes. These efforts have spanned different representation spaces, including the UNet bottleneck layer (Kwon, Jeong, and Uh 2022; Park et al. 2023), text embedding space (Gandikota et al. 2024; Baumann et al. 2025; Hertz et al. 2022), noise space (Dalva and Yanardag 2024), weight space (Dravid et al. 2024; Gandikota et al. 2024), and representation in GAN-based architecture (Li, Hou, and Loy 2024; Tov et al. 2021). However, these methods often suffer from semantic entanglement (Kwon, Jeong, and Uh 2022; Park et al. 2023), lack clear attribute correspondence (Dalva and Yanardag 2024), or rely on the addition of phrases with limited continuity of attribute manipulation (Hertz et al. 2022). Some other studies (Li, Hou, and Loy 2024; Tov et al. 2021) achieve precise control, but their methods are designed for the GAN spaces, restricting the generalization to the powerful diffusion models. As a result, these methods fail to achieve both fine-grained and continuous attribute manipulation within a single framework.
@@
Recent studies (Dravid et al. 2024; Gandikota et al. 2024; Baumann et al. 2025) have made efforts in achieving continuous attribute manipulation (e.g., gradually intensifying one’s smile or progressively increasing the degree of aging). Some methods  (Dravid et al. 2024; Gandikota et al. 2024) rely on domain-specific LoRA training to achieve latent space disentanglement. Others (Baumann et al. 2025) introduce attribute-wise training, which enables continuous attribute manipulation but requires attribute-wise supervision and heavily relies on labeled or paired data. As a result, these methods typically require separate training for each attribute and follow a One-for-One design paradigm. In their pipeline, each attribute is handled in isolation via a dedicated slider module or a specific optimized representation direction, as shown in Figure 2(1). When manipulating multiple attributes, this One-for-One paradigm requires multiple training runs, leading to significant time and computational costs and limiting their scalability and applicability in real-world scenarios.
@@
Facing the limitations of the One-for-One paradigm, we propose an All-in-One Slider framework, as illustrated in Figure 2(2), which uses a single lightweight module to achieve flexible control over multiple attributes. In order to achieve generalized control of multiple attributes within a single module, a key challenge lies in disentangling these attributes. Once disentangled, the intensity of a target attribute can be easily controlled without affecting others, therefore enabling generalized and precise attribute manipulation. Inspired by the recent studies of sparse autoencoder (SAE) in LLMs (Karvonen et al. 2024; Huben et al. 2023), we adopt their idea of “break it down to build it up” and propose an Attribute Sparse Autoencoder, which is implemented on the text embeddings used in the image generation process. It first encodes each embedding into many fundamental components in a unified latent space for diverse attributes. By controlling the activation of these components, the desired attributes are then semantically created through decoding the activated ones. Specifically, we collect a set of commonly used textual prompts in portrait generation as our training dataset and encode them into the text embeddings. Subsequently, the Attribute Sparse Autoencoder is trained to perform auto-encoding on these embeddings and activate as sparsely as possible, in which the sparsity is the key to attribute disentanglement. During inference, we can control the activation of corresponding components in the latent space. By directly manipulating the coefficient of activated components, it naturally supports continuous and fine-grained control over desired attributes.
Through operating in this sparsely activated latent space of attributes, our All-in-One Slider can achieve decoupled control of multiple attributes through a single module, without the need for any specific training for each attribute. Moreover, as a unified attribute manipulation framework, our All-in-One Slider shows scalability and generalization to unseen attributes in T2I generation. Both qualitative and quantitative experiments are conducted to demonstrate the effectiveness of our All-in-One Slider for continuous and fine-grained manipulation of attributes.
@@
Related Work
Attribute Manipulation in T2I Diffusion Models
Controllability has become a central goal in T2I diffusion models, enabling users to steer generation toward desired contents or styles. Early approaches, such as ControlNet (Zhang, Rao, and Agrawala 2023; Mou et al. 2024), introduce spatial guidance by conditioning on structural cues like edges, poses, or depth maps. These methods are effective at preserving geometric alignment but are inherently limited when it comes to controlling high-level semantic attributes—such as age, expression, or makeup—which are not explicitly represented in these inputs.
To achieve fine-grained semantic control beyond spatial guidance, researchers explored semantic manipulation in latent spaces (e.g., noise or $\mathcal{H}$-space). Methods like SEGA (Brack et al. 2023) and Boundary Guidance (Zhu et al. 2023) enable editing without training, but lack fine-grained continuity or compositional control over semantic attributes. Later works shifted toward manipulating the text embedding space to achieve more direct semantic control. Prompt-based methods (Hertz et al. 2022; Brooks, Holynski, and Efros 2023) intervene through text prompts and attention maps, but often suffer from discrete or coarse edits. Meanwhile, editing methods like Imagic (Kawar et al. 2023) and ITI-Gen (Zhang et al. 2023) focus on aligning generation with user-provided images or preferences, but rely on image-text pairs and lack scalability for general attribute manipulation. Recent approaches, like ConceptSlider (Gandikota et al. 2024) and AttributeControl (Baumann et al. 2025), seek to enable continuous control by learning attribute vectors with paired data. However, they still require training a dedicated slider for each attribute, which limits scalability to unseen attributes.
@@
Sparse Autoencoders (SAEs)
Sparse autoencoders (Ng et al. 2011; Olshausen and Field 1997; Bengio, Courville, and Vincent 2013) have been proven effective in learning interpretable representations in unsupervised settings (Coates, Ng, and Lee 2011). Among them, K-Sparse Autoencoders (K-SAE) (Makhzani and Frey 2013) enforce sparsity by allowing only the top-$k$ activations to activate, promoting the emergence of semantically meaningful units.
Some studies in the context of large language models (Gao et al. 2024; Cunningham et al. 2023) introduce SAEs in capturing discrete semantic concepts. Recent studies leverage the SAEs to interpret and analyze the internal mechanisms of diffusion models. SAeUron (Cywiński and Deja 2025) uses SAEs for concept unlearning by suppressing undesired features across diffusion steps. Diffusion Lens (Toker et al. 2024) decodes intermediate text encoder activations to reveal semantic representations. Unpacking SDXL Turbo (Surkov et al. 2025) decomposes transformer blocks to expose interpretable, causal factors across style, structure, and semantics. However, these works primarily focus on model interpretability and causal tracing. In this work, we extend K-SAE into a controllable face attribute manipulation framework that serves as a lightweight and scalable solution, enabling precise attribute control without retraining the base model.
Method
@@
Preliminary: Diffusion-based T2I Generation
With the advent of DDPM (Ho, Jain, and Abbeel 2020), diffusion models (Saharia et al. 2022; Nichol et al. 2021) have become the dominant technology in visual generation, especially for many state-of-the-art T2I generation systems. Recent T2I diffusion models typically follow the latent generation process (Rombach et al. 2022) and consist of three main components: a VAE (Kingma and Welling 2013), a text encoder, and a diffusion generator. The VAE compresses images into a latent space and decodes the generated latent back to pixel space. The text encoder, such as CLIP (Radford et al. 2021), encodes textual prompts into embeddings. The diffusion generator then takes a noisy latent as input and captures prompt information through cross-attention with the text embeddings, iteratively denoising the noisy latent towards a final output, enabling text-to-image generation.
@@
Attribute Disentanglement Training
To achieve encoding and memorization of as many attributes as possible with limited parameters, we introduce the Attribute Sparse Autoencoder, which contains an encoder and a decoder. For training, we collect numerous text embeddings from the text encoder of a pretrained diffusion model. The encoder part of the Attribute Sparse Autoencoder maps these embeddings into a high-dimensional latent space (referred to as Attlatentspace in the following). Once trained, this Attlatentspace can serve as a unified attribute space, where attributes with different semantics sparsely activate a minimal and distinct set of neurons, thereby enabling disentangled modeling of diverse attributes. The decoder part then reconstructs the original embeddings from the latent representations. To encourage sparsity and disentanglement in the Attlatentspace, we draw inspiration from SAEs (Makhzani and Frey 2013; Gao et al. 2024; Karvonen et al. 2024) and apply a top-$k$ activation strategy to sparsely activate the attribute latent space. The overall training procedure consists of three major steps: embedding extraction, sparse encoding, and sparse reconstruction, as detailed below.
Embedding Extraction
The text encoder of a pretrained diffusion model (e.g., SDXL) consists of a stack of transformer blocks $\{B_{1},\ldots,B_{L}\}$. We extract the hidden state $x$ from an intermediate block $B_{l}$ within this stack. Specifically, we extract features from both SDXL text encoders, concatenate them to form the final joint representation $x$, and feed it into the Attribute Sparse Autoencoder.
Sparse Encoding
We implement sparsity using a top-$k$ selection mechanism. The encoder ENC maps the extracted embedding $x$ into the Attlatentspace $z_{\text{ALS}}$ via a linear transformation, ReLU activation, and top-$k$ selection:
@@
Here, retaining only the top-$k$ neurons with the highest activation yields a compact and disentangled representation, where each attribute is associated with a distinct subset of neurons. This sparsity facilitates precise attribute manipulation during generation. The decoder then reconstructs $\hat{x}$ as:
@@
Sparse Reconstruction
We train Attribute Sparse Autoencoder to minimize the reconstruction loss:
@@
However, a known limitation of sparse autoencoders is the prevalence of dead neurons—latent units that often remain inactive throughout training. These unused neurons reduce the expressiveness of the learned space and hinder future composability. To address this, we adopt an auxiliary mechanism inspired by (Gao et al. 2024). At each training step, we compute the residual $r=x-\hat{x}$ and identify a set of $k_{\text{aux}}$ least-activated neurons. The corresponding activations, denoted as $\hat{z}_{\text{ALS}}$, are used to reconstruct the residual:
@@
This loss encourages even initially inactive neurons to explain parts of the residual, ensuring broader semantic coverage in the latent space. After training, each dimension in the sparse latent space corresponds to a specific, disentangled facial attribute. This latent space serves as the basis for attribute manipulation, as detailed in the next section.
Attribute Manipulation
@@
Once the Attribute Sparse Autoencoder is trained, it can be employed as an “Attribute Manipulation Slider”, enabling control over specific attributes within the unified Attribute Latent Space of the text encoder. Given an attribute $A$ that we expect to manipulate, we extract its corresponding text embedding $x_{A}$ and feed it to the trained Attribute Sparse Autoencoder to identify the latent direction associated with the attribute. We then modify the input prompt embedding $x$ by adding this direction scaled by a scalar $\lambda$, as follows:
$x_{\text{manipulated}}=x+W_{\text{dec}}(\lambda\times\text{ENC}(x_{A})),$ (6)
@@
where $\lambda$ controls the strength of the steering. This modification of the embedding $x_{\text{manipulated}}$ is then used to condition the generation process, enabling precise manipulation of specific attributes without disrupting the overall quality or structure of the generated image. Importantly, this manipulation operates entirely within the learned sparse latent space, supporting zero-shot, composable editing across a wide range of diverse attributes. This unified mechanism underpins the scalable and flexible attribute control explored.
Experiments
Experimental Setup
Implementation Details

### v2

@@ Abstract
Text-to-image (T2I) diffusion models have made significant strides in generating high-quality images. However, progressively manipulating certain attributes of generated images to meet the desired user expectations remains challenging, particularly for content with rich details, such as human faces. Some studies have attempted to address this by training slider modules. However, they follow a One-for-One manner, where an independent slider is trained for each attribute, requiring additional training whenever a new attribute is introduced. This not only results in parameter redundancy accumulated by sliders but also restricts the flexibility of practical applications and the scalability of attribute manipulation. To address this issue, we introduce the All-in-One Slider, a lightweight module that decomposes the text embedding space into sparse, semantically meaningful attribute directions. Once trained, it functions as a general-purpose slider, enabling interpretable and fine-grained continuous control over various attributes. Moreover, by recombining the learned directions, the All-in-One Slider supports the composition of multiple attributes and zero-shot manipulation of unseen attributes (e.g., races and celebrities). Extensive experiments demonstrate that our method enables accurate and scalable attribute manipulation, achieving notable improvements compared to previous methods. Furthermore, our method can be extended to integrate with the inversion framework to perform attribute manipulation on real images, broadening its applicability to various real-world scenarios. The code is available on our project page.
1 Introduction
@@
With the widespread emergence of diffusion models [44, 34], text-to-image (T2I) generation has advanced significantly in recent years. Despite their impressive performance, a major challenge remains: providing users with continuous and fine-grained control functions over specific image attributes, which is essential for users to adjust generated images to obtain the desired results. As illustrated by the example in Figure 1(1), traditional prompt engineering approaches, such as simply appending phrases like “with a big smile” to the base prompt, “A photo of a gentle Thai man with layered hair, in a warm-toned studio”, often result in coarse-grained and rigid manipulations. Such textual variations not only prevent users from specifying the intensity of the smile but also inadvertently affect many unrelated attributes, such as the subject’s hairstyle, pose, and identity, leading to suboptimal outcomes.
@@
Figure 1: Our All-in-One Slider shows advantages in: (1) Fine-grained and continuous control over desired attribute, without affecting other attributes (e.g., subject identity and appearance). (2) Combination of multiple facial attributes (e.g., smile and age) for consistent and conflict-free transformations. (3) Zero-shot generalization to unseen attributes, without multiple and cumbersome training processes.
@@
To enable attribute manipulation, early studies focused on exploring the representation properties of generative models and identifying the semantic directions corresponding to attributes. These efforts have spanned different representation spaces, including the UNet bottleneck layer [25, 33], text embedding space [13, 1, 17], noise space [9], weight space [12, 13], representation in GAN-based architecture [27, 42], and more recently, rectified flow trajectories [10, 37]. However, these methods often suffer from semantic entanglement [25, 33], lack clear attribute correspondence [9], or rely on the addition of phrases with limited continuity of attribute manipulation [17]. Some other studies [27, 42] achieve precise control, but their methods are designed for the GAN spaces, restricting the generalization to the powerful diffusion models. As a result, these methods fail to achieve both fine-grained and continuous attribute manipulation within a single framework.
@@
Recent studies [12, 13, 1, 6] have made efforts in achieving continuous attribute manipulation (e.g., gradually intensifying one’s smile or progressively increasing the degree of aging). Some methods  [12, 13, 6] rely on domain-specific LoRA or adapter training to achieve latent space disentanglement. Others [1] introduce attribute-wise training, which enables continuous attribute manipulation but requires attribute-wise supervision and heavily relies on labeled or paired data. As a result, these methods typically require separate training for each attribute and follow a One-for-One design paradigm. In their pipeline, each attribute is handled in isolation via a dedicated slider module or a specific optimized representation direction, as shown in Figure 2(1). When manipulating multiple attributes, this One-for-One paradigm requires multiple training runs, leading to significant time and computational costs and limiting their scalability and applicability in real-world scenarios.
@@
Facing the limitations of the One-for-One paradigm, we propose an All-in-One Slider framework, as illustrated in Figure 2(2), which uses a single lightweight module to achieve flexible control over multiple attributes. In order to achieve generalized control of multiple attributes within a single module, a key challenge lies in disentangling these attributes. Once disentangled, the intensity of a target attribute can be easily controlled without affecting others, therefore enabling generalized and precise attribute manipulation. Inspired by the recent studies of sparse autoencoder in LLMs [22, 21, 47], we adopt the idea of “break it down to build it up” and propose an Attribute Sparse Autoencoder, which is implemented on the text embeddings used in the image generation process. It first maps the text embedding of the attribute into a unified latent space (high-dimensional space), where individual semantic components are represented by sparse neuron activations. By selectively activating these neurons, the module can reconstruct and control specific attributes in a disentangled manner through decoding. Specifically, we collect common textual prompts in portrait generation as our training dataset and encode them into the text embeddings. Subsequently, the Attribute Sparse Autoencoder is trained to perform auto-encoding on these embeddings and activate as sparsely as possible, in which the sparsity is the key to attribute disentanglement. Through this training, the unified latent space forms a compact, compositional dictionary of semantic components that can be flexibly recombined to represent new or unseen attributes without any retraining. During inference, we can control the activation of corresponding components in the latent space. By directly manipulating the coefficient of activated components, it naturally supports continuous and fine-grained control over desired attributes.
Through operating in this sparsely activated latent space of attributes, our All-in-One Slider can achieve decoupled control of multiple attributes through a single module, without the need for any specific training for each attribute. Benefiting from the compositional nature of the unified latent space, our All-in-One Slider generalizes to a wide range of attributes and enables continuous zero-shot manipulation of previously unseen ones. Both qualitative and quantitative experiments are conducted to demonstrate the effectiveness of our All-in-One Slider for continuous and fine-grained manipulation of attributes.
@@
2 Related Work
Attribute Manipulation in T2I Diffusion Models. Achieving precise controllability has become a central goal in T2I diffusion models, enabling users to steer generation toward desired contents or styles. Early approaches, such as ControlNet [46, 29], introduce spatial guidance by conditioning on structural cues like edges, poses, or depth maps. These methods are effective at preserving geometric alignment but are inherently limited in controlling high-level semantic attributes, e.g., age, expression, or makeup, which are not explicitly represented in these inputs.
To achieve fine-grained semantic control beyond spatial guidance, researchers explored semantic manipulation in latent spaces (e.g., noise or $\mathcal{H}$-space). Methods like SEGA [4] and Boundary Guidance [48] enable editing without training, but lack fine-grained continuity or compositional control over semantic attributes. Later works shifted toward manipulating the text embedding space to achieve more direct semantic control. Prompt-based methods [17, 5] intervene through text prompts and attention maps, but often suffer from discrete or coarse edits. Meanwhile, editing methods like Imagic [23] and ITI-Gen [45] focus on aligning generation with user-provided images or preferences, but rely on image-text pairs and lack scalability for general attribute manipulation. Recent approaches, like ConceptSlider [13] and AttributeControl [1] (AttControl), seek to enable continuous control by learning attribute vectors. However, for each attribute, they all require training a dedicated slider from lots of pairwise data, which leads to parameter redundancy accumulated by sliders and limits flexibility and scalability in practical application.
@@ 2 Related Work
Sparse Autoencoders (SAEs). Sparse autoencoders [30, 32, 2] have been proven effective in learning interpretable representations in unsupervised settings [7]. Among them, K-Sparse Autoencoders (K-SAE) [28] enforce sparsity by allowing only the top-$k$ activations to activate, promoting the emergence of semantically meaningful units. Some studies in the context of large language models [14, 21] introduce SAEs in capturing discrete semantic concepts. Recent studies leverage the SAEs to interpret and analyze the internal mechanisms of diffusion models. SAeUron [8] uses SAEs for concept unlearning by suppressing undesired features across diffusion steps. Diffusion Lens [41] decodes intermediate text encoder activations to reveal semantic representations. Unpacking SDXL Turbo [40] decomposes transformer blocks to expose interpretable, causal factors across style, structure, and semantics. However, these works primarily focus on model interpretability and causal tracing. In this work, we extend K-SAE into a controllable face attribute manipulation framework that serves as a lightweight and scalable solution, enabling precise attribute control without retraining the base model.
@@
3 Method
@@
3.1 Preliminary: Diffusion-based T2I Generation
With the advent of DDPM [20], diffusion models [38, 31] have become the dominant technology in visual generation, especially for many state-of-the-art T2I generation systems. Recent T2I diffusion models typically follow the latent generation process [36] and consist of three main components: a Variational Auto-Encoder (VAE) [24], a text encoder, and a diffusion generator. The VAE compresses images into a latent space and decodes the generated latent back to pixel space. The text encoder, such as CLIP [35], encodes textual prompts into embeddings. The diffusion generator then takes a noisy latent as input and captures prompt information through cross-attention with the text embeddings, iteratively denoising the noisy latent towards a final output, enabling text-to-image generation.
@@ 3.1 Preliminary: Diffusion-based T2I Generation
3.2 Attribute Disentanglement Training
To achieve encoding and memorization of as many attributes as possible with limited parameters, we introduce the Attribute Sparse Autoencoder, which contains an encoder and a decoder. For training, we collect numerous text embeddings from the text encoder of a pretrained diffusion model. The encoder part of the Attribute Sparse Autoencoder maps these embeddings into a high-dimensional latent space (referred to as Attlatentspace in the following). Once trained, this Attlatentspace can serve as a unified attribute space, where attributes with different semantics sparsely activate a minimal and distinct set of neurons, thereby enabling disentangled modeling of diverse attributes. The decoder part then reconstructs the original embeddings from the latent representations. To encourage sparsity and disentanglement in the Attlatentspace, we draw inspiration from SAEs [28, 14, 22] and apply a top-$k$ activation strategy to sparsely activate the attribute latent space. The overall training procedure consists of three major steps: embedding extraction, sparse encoding, and sparse reconstruction, as detailed below.
Embedding Extraction. The text encoder of a pretrained diffusion model (e.g., SDXL) consists of a stack of transformer blocks $\{B_{1},\ldots,B_{L}\}$. We extract the hidden state $x$ from an intermediate block $B_{l}$ within this stack. Specifically, we extract features from the dual text encoders of SDXL, concatenate them to form the final joint representation $x$, and feed it into the Attribute Sparse Autoencoder.
Sparse Encoding. We implement sparsity using a top-$k$ selection mechanism. The encoder ENC maps the extracted embedding $x$ into the Attlatentspace $z_{\text{ALS}}$ via a linear transformation, ReLU activation, and top-$k$ selection:
@@ 3.2 Attribute Disentanglement Training
Here, retaining only the top-$k$ neurons with the highest activation yields a compact and disentangled representation, where each attribute is associated with a distinct subset of neurons. The sparsity is applied to the individual channels of the text embedding, enabling the selective activation of features that are most pertinent to the semantics of specific attributes in the high-dimensional embedding space (i.e., Attlatentspace). This channel-wise sparsity thereby promotes attribute disentanglement, allowing the model to isolate fine-grained semantic signals while suppressing irrelevant noise. The decoder then reconstructs $\hat{x}$ as:
@@
Sparse Reconstruction. We train Attribute Sparse Autoencoder to minimize the reconstruction loss:
@@
However, a known limitation of sparse autoencoders is the prevalence of dead neurons—latent units that often remain inactive throughout training. These unused neurons reduce the expressiveness of the learned space and hinder future composability. To address this, we adopt an auxiliary mechanism inspired by [14]. At each training step, we compute the residual $r=x-\hat{x}$ and identify a set of $k_{\text{aux}}$ least-activated neurons. The corresponding activations, denoted as $\hat{z}_{\text{ALS}}$, are used to reconstruct the residual:
@@
This loss encourages even initially inactive neurons to explain parts of the residual, ensuring broader semantic coverage in the latent space. After training, latent units in the sparse latent space correspond to specific, disentangled facial attributes. This latent space serves as the basis for attribute manipulation, as detailed in the next section.
3.3 Attribute Manipulation
@@
Once the Attribute Sparse Autoencoder is trained, it can be employed as an “Attribute Manipulation Slider”, enabling control over specific attributes within the unified Attribute Latent Space of the text encoder. Given an attribute $A$ that we expect to manipulate, we extract its corresponding text embedding $x_{\text{A}}$ and feed it to the trained Attribute Sparse Autoencoder to identify the latent direction associated with the attribute. We then modify the input prompt embedding $x$ by adding this direction scaled by a scalar $\lambda$, as follows:
$x_{\text{manipulated}}=x+W_{\text{dec}}(\lambda\times\text{ENC}(x_{\text{A}})),$ (6)
@@
where $\lambda$ controls the strength of the steering. This modification of the embedding $x_{\text{manipulated}}$ is then used to condition the generation process, enabling precise manipulation of specific attributes without disrupting the overall quality or structure of the generated image. Importantly, this manipulation operates entirely within the learned sparse latent space, supporting zero-shot, composable editing across a wide variety of attributes. This unified mechanism underpins the scalable and flexible attribute control explored.
4 Experiments
4.1 Experimental Setup